OpenStudy (anonymous):

Factor the algebraic expression below in terms of a single trigonometric function. csc^2x - 1

OpenStudy (sidsiddhartha):

cscx=1/sinx csc^x-1=1/sin^2x-1=(1-sin^2x)/sin^2x=cos^2x/sin^2x=cot^2x

OpenStudy (anonymous):

...Wow, I'm totally lost. -shakes head.- Okay, let me try to trace over what you did. First off. \[\csc ^{2}x - 1 \] Then. You do.. \[\csc x-1 =1 \div \sin^2-1 \] But why and how?

OpenStudy (sidsiddhartha):

sorry it's csc^2x-1=(1/sin^2x)-1

OpenStudy (anonymous):

Okay, so since 1/sin^2x is the reciprocal function. Thats how you got it, simply switching around. But why would you divide it by sin and where does the -1 go to?

OpenStudy (sidsiddhartha):